OpenStudy (zeesbrat3):

If f(x) = ∣(x2 − 8)∣, how many numbers in the interval 0 ≤ x ≤ 2.5 satisfy the conclusion of the mean value theorem?

OpenStudy (zeesbrat3):

Please help

OpenStudy (amoodarya):

if u plot it , it will be simpler to do

OpenStudy (amoodarya):

\[|x^2-8|=|(x-2\sqrt{2})(x+2\sqrt{2})|\]

OpenStudy (amoodarya):